# Reconstruction of QCD first-order phase transition from neutron star measurements

### v1

@@ III.3 Marginal posterior distribution of phase transition parameters
With the detail of computation given in Appendix B.4, we compute $P(\xi_{\mathrm{PT}},\Delta\varepsilon)$ using the same setup as in the preceding subsection and consider two different levels of uncertainties: a more realistic level with $\Delta_{M}=0.01~{}M_{\odot}$, $\Delta_{R}=0.1~{}\mathrm{km}$, and a more optimistic level $\Delta_{M}=0.1~{}M_{\odot}$, $\Delta_{R}=1.0~{}\mathrm{km}$. Results are shown in Fig. 4, where the orange stars represent the true PT parameters listed in the preceding subsection, whereas the white crosses indicate those maximizing the corresponding marginal posteriors. Note that they do not necessarily coincide with the MAP values listed above since other parameters have been integrated.
@@ IV Summary and Discussions
With the newly developed algorithm, we employed a general, unbiased parametrization of the EoS realized by neural networks, and demonstrated its ability in reconstruction of the EoS from NS measurements with and without a first-order PT. We further explored the ability of this method to reconstruct PT parameters from noisy mass and radius measurements. Based on the settings in the mock observations, we found that measurements with the state-of-the-art precisions is able to reveal a strong PT with magnitude $\Delta\varepsilon\sim 100~{}\mathrm{MeV}/\mathrm{fm}^{3}$.
While the above statement can be treated as a semi-quantitative statement of the order-of-magnitude, we admit that the exact value is unavoidably model dependent — it could rely on the based EoS, number and distribution of the NS measurements, and the critical pressure at which the PT occurs. Note that the quantitative analysis in Sec. III.3 has already taken $\sim 8\times 10^{5}\,\mathrm{cpu}\cdot\mathrm{hours}$; the model dependence shall be analyzed in a neater way in the future.
Meanwhile, based on the sensitivity analysis the tidal deformability is more sensitive to the EoS than the mass and radius, and we expect that possible future measurements of the $M$-$R$-$\Lambda$ relation shall be able to better constrain the EoS and detect the possible QCD phase transition. It is worth noting that the linear response analysis also applies to other observables of NSs, such as the moment of inertia and quadrupole moment. Such analyses could shed lights on the understanding of the approximate universal relations among these quantities.
@@ Acknowledgments
The authors thank Zoey Zhiyuan Dong, Christian Drischler, Philippe Landry, Lap-Ming Lin, Shriya Soma, Andrew Steiner, and Boyang Sun for helpful discussions. We gratefully acknowledge the DEEP-IN working group at RIKEN-iTHEMS for support in the preparation of this paper. This work is supported by Tsinghua University under grant No. 043-04200500123, No. 043-531205006, and No. 043-533305009 (S.S.), Guangdong Major Project of Basic and Applied Basic Research under grant No. 2020B030103000 (R.L.), Startup Funds from the T.D. Lee Institute and Shanghai Jiao Tong University (S.H.), NSF under grant No. PHY 21-16686 (Z.L.), the RIKEN TRIP initiative (RIKEN Quantum) and JST-BOOST grant (L.W.), the CUHK-Shenzhen university development fund under grant No. UDF01003041 and No. UDF03003041, and Shenzhen Peacock fund under No. 2023TC0179 (K.Z.).
@@ Appendix A Numerical Validation of the Linear Response Analysis
In this Appendix, we numerically validate the derivatives computed in Sec. II. We start from the SFHo EoS, compute the NS mass, radius, and tidal deformability for various central pressures ({$P_{c}$}). The obtained lists are denoted as $\{M\}$, $\{R\}$, $\{\Lambda\}$. To compute the numerical derivatives with respect to the central pressure, we repeat the procedure for a list of slightly different central pressures, $\{P_{c}\times(1+\delta\ln\frac{P_{c}}{P_{\text{bnd}}})\}$, and the corresponding observables are $\{\widetilde{M}\}$, $\{\widetilde{R}\}$, $\{\widetilde{\Lambda}\}$. Numerical derivatives are then computed as $\{\frac{\widetilde{M}-M}{\delta\ln(P_{c}/P_{\text{bnd}})}\}$, $\{\frac{\widetilde{R}-R}{\delta\ln(P_{c}/P_{\text{bnd}})}\}$, and $\{\frac{\widetilde{\Lambda}-\Lambda}{\Lambda\delta\ln(P_{c}/P_{\text{bnd}})}\}$, which are compared to the results computed using Eqs. (34, 3538) and shown in the top panels of Fig. 5. $\delta\ln\frac{P_{c}}{P_{\text{bnd}}}$ is taken to be $\sim 0.04$ in numerical calculations.
@@
Numerical derivatives are then computed as $\{\frac{\overline{M}-M}{\Delta\varepsilon}\}$, $\{\frac{\overline{R}-R}{\Delta\varepsilon}\}$, and $\{\frac{\overline{\Lambda}-\Lambda}{\Delta\varepsilon}\}$, which are compared to the results computed using Eqs. (2628, 3538) and shown in the bottom panels of Fig. 5. In numerical calculations, we take $P_{\mathrm{PT}}=60~{}\text{MeV/fm}^{3}$ and $\Delta\varepsilon_{\mathrm{PT}}=1~{}\text{MeV/fm}^{3}$.

### v2

@@ III.3 Marginal posterior distribution of phase transition parameters
With the detail of computation given in Appendix B.4, we compute $P(\xi_{\mathrm{PT}},\Delta\varepsilon)$ using the same setup as in the preceding subsection and consider two different levels of uncertainties: a more optimistic level with $\Delta_{M}=0.01~{}M_{\odot}$, $\Delta_{R}=0.1~{}\mathrm{km}$, and a more realistic level $\Delta_{M}=0.1~{}M_{\odot}$, $\Delta_{R}=1.0~{}\mathrm{km}$. Results are shown in Fig. 5, where the orange stars represent the true PT parameters listed in the preceding subsection, whereas the white crosses indicate those maximizing the corresponding marginal posteriors. Note that they do not necessarily coincide with the MAP values listed above since other parameters have been integrated.
@@ IV Summary and Discussions
With the newly developed algorithm, we employed a general, unbiased parametrization of the EoS realized by neural networks, and demonstrated its ability in reconstruction of the EoS from NS measurements with and without a first-order PT. We further explored the ability of this method to reconstruct PT parameters from noisy mass and radius measurements. Based on the settings in the mock observations, we can make an order-of-magnitude estimate that measurements with the state-of-the-art precisions shall be able to reveal a strong PT with $\Delta\varepsilon\sim\mathcal{O}(10^{2})~{}\mathrm{MeV}/\mathrm{fm}^{3}$. We admit that the exact value is unavoidably model dependent — it could rely on the based EoS, the number and distribution of the NS measurements, and the critical pressure at which the PT occurs. Note that the quantitative analysis in Sec. III.3 has already taken $\sim 8\times 10^{5}\,\mathrm{cpu}\cdot\mathrm{hours}$; the model dependence shall be analyzed in a neater way in the future.
Meanwhile, according to the sensitivity analysis, the tidal deformability turns out to be more sensitive to the EoS than the NS mass and radius, and we expect that possible future measurements of the $M$-$R$-$\Lambda$ relation shall be able to better constrain the EoS and detect the possible QCD phase transition. Explorations of the PT reconstruction using potential tidal deformability measurements of NSs will be performed in the future. It is worth noting that the linear response analysis also applies to other observables of NSs, such as the moment of inertia and quadrupole moment. Such analyses could shed lights on the understanding of the approximate universal relations among these quantities.
@@ Acknowledgments
The authors thank Zoey Zhiyuan Dong, Christian Drischler, Philippe Landry, Lap-Ming Lin, Shriya Soma, Andrew Steiner, and Boyang Sun for helpful discussions. We gratefully acknowledge the DEEP-IN working group at RIKEN-iTHEMS for support in the preparation of this paper. This work is supported by the National Key Research and Development Program of China under Contract No. 2024YFA1610700 (S.S.), Tsinghua University under grant No. 043-04200500123, No. 043-531205006, and No. 043-533305009 (S.S.), Guangdong Major Project of Basic and Applied Basic Research under grant No. 2020B030103000 (R.L.), Startup Funds from the T.D. Lee Institute and Shanghai Jiao Tong University (S.H.), NSF under grant No. PHY 21-16686 (Z.L.), the RIKEN TRIP initiative (RIKEN Quantum) and JST-BOOST grant (L.W.), the CUHK-Shenzhen university development fund under grant No. UDF01003041 and No. UDF03003041, and Shenzhen Peacock fund under No. 2023TC0179 (K.Z.).
Data Availability
The data that support the findings of this article are openly available [105].
@@ Appendix A Numerical Validation of the Linear Response Analysis
In this Appendix, we numerically validate the derivatives computed in Sec. II. We start from the SFHo EoS, compute the NS mass, radius, and tidal deformability for various central pressures ({$P_{c}$}). The obtained lists are denoted as $\{M\}$, $\{R\}$, $\{\Lambda\}$. To compute the numerical derivatives with respect to the central pressure, we repeat the procedure for a list of slightly different central pressures, $\{P_{c}\times(1+\delta\ln\frac{P_{c}}{P_{\text{bnd}}})\}$, and the corresponding observables are $\{\widetilde{M}\}$, $\{\widetilde{R}\}$, $\{\widetilde{\Lambda}\}$. Numerical derivatives are then computed as $\{\frac{\widetilde{M}-M}{\delta\ln(P_{c}/P_{\text{bnd}})}\}$, $\{\frac{\widetilde{R}-R}{\delta\ln(P_{c}/P_{\text{bnd}})}\}$, and $\{\frac{\widetilde{\Lambda}-\Lambda}{\Lambda\delta\ln(P_{c}/P_{\text{bnd}})}\}$, which are compared to the results computed using Eqs. (32, 3336) and shown in the top panels of Fig. 6. $\delta\ln\frac{P_{c}}{P_{\text{bnd}}}$ is taken to be $\sim 0.04$ in numerical calculations.
@@
Numerical derivatives are then computed as $\{\frac{\overline{M}-M}{\Delta\varepsilon}\}$, $\{\frac{\overline{R}-R}{\Delta\varepsilon}\}$, and $\{\frac{\overline{\Lambda}-\Lambda}{\Delta\varepsilon}\}$, which are compared to the results computed using Eqs. (2426, 3336) and shown in the bottom panels of Fig. 6. In numerical calculations, we take $P_{\mathrm{PT}}=60~{}\text{MeV/fm}^{3}$ and $\Delta\varepsilon_{\mathrm{PT}}=1~{}\text{MeV/fm}^{3}$.